A water bottle company produces bottles with a net volume 400ml each. It is known that the standard deviation of the volume is 12ml. A sample of 50 bottles is tested.


If the sample yields mean volume of 420ml, construct a hypothesis test at the \alpha =0.01 level of significance for the null hypotheses H_(0):\mu =400ml against the alternative hypothesis H_(A):\mu !=400ml. State your conclusion and give a reason for your answer.

1
Expert's answer
2021-11-30T09:32:16-0500

The hypotheses tested are,

"H_0:\mu= 400"

"Against"

"H_A:\mu\not=400"

We are given the following,

"\bar{x}=420,\space \sigma=12,\space n=50"

We will apply the standard normal distribution to perform this hypothesis test as follows.

"Z_c=(\bar{x}-\mu)/(\sigma/\sqrt{n})=(420-400)/(12/\sqrt{50})=20/1.6971=11.785113"

"Z_c" is compared with the standard normal table value at "\alpha=0.01" given as, "Z_{0.01\/2}=Z_{0.005}=2.575" and the null hypothesis is rejected if "Z_c\gt Z_{0.005}."

Since "Z_c=11.785113\gt Z_{0.005}=2.575", we reject the null hypothesis and we conclude that there is no sufficient evidence to show that the average net volume for bottles is 400 ml at 1% level of significance.